OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. A 2 kg laptop sits on the floor near a 4 kg jar of pennies. If the force of gravity between them is 3.42 10-10 N, how far apart are they? A. 2.1 m B. 1.6 m C. 1.2 m D. 1.8 m

OpenStudy (theeric):

For reference, you can still look at http://openstudy.com/study#/updates/52c77b0fe4b0095f1003fd66 So, you're looking at the force of gravity. And it looks like you want the Universal Law of Gravitation. It's a relationship that can be written as an equation. Do you know what that equation is?

OpenStudy (anonymous):

is the equation still F=G=m1*m2/r2?

OpenStudy (theeric):

F=G\(\large\text *\)m1*m2/r\(\large\text ^\)2

OpenStudy (theeric):

Do you see the differences? And that is the equation :)

OpenStudy (anonymous):

so what number would be r^2? i have it up to there i think. so far it's 3.42*2*4/? @theEric

OpenStudy (theeric):

F=G*m1*m2/ r^2 F=3.42*10^-10 G=6.672*10^-11 m1=2 m2=4 r=?? unknown ??

OpenStudy (theeric):

I should've spaced that out better... The equation is at the top, and the given values are below it.

OpenStudy (theeric):

G=6.674*10^-11, sorry!

OpenStudy (anonymous):

so if r is unknown, that must be what im trying to find. So wouldnt the equation be switched around to where it says r=then the equation? @theEric

OpenStudy (theeric):

Yep!! :) Now you're getting there! :)

OpenStudy (theeric):

Give it a shot!

OpenStudy (anonymous):

I honestly have no idea. would you just switch r & f? so it's r=g*m1*m2/f?

OpenStudy (anonymous):

@theEric

OpenStudy (theeric):

Yes you would!!! Now you can use a calculator, and substitute in the values I listed above!

OpenStudy (theeric):

Well wait...

OpenStudy (theeric):

that is r^2! r squared! \(r^2\) So make sure you see that. You can take the square root of both sides to finish solving for r!

OpenStudy (theeric):

Then substitute.

OpenStudy (anonymous):

i got 15.6 what did i do wrong? -.- @theEric

OpenStudy (theeric):

Two things! One, you didn't remember to use the times 10^-11 and 10^-10. Two, you didn't take the square root of it all. Don't worry, you're getting there!

OpenStudy (anonymous):

I got 1.2 is that correct? @theeric

OpenStudy (theeric):

Final answer? I agree. Congratulations! :)